OpenStudy (katecc379):

Given the arithmetic sequence an = 3 + 2(n − 1), what is the domain for n?

OpenStudy (katecc379):

@retirEEd

OpenStudy (katecc379):

@kidrah69

OpenStudy (katecc379):

is it all integers greater than or equal to 1

OpenStudy (retireed):

Why can't n be a negative or zero?

OpenStudy (katecc379):

cause domains for sequences are always positive numbers

OpenStudy (katecc379):

right?

OpenStudy (retireed):

I really don't know that to be true, so I will have to go to Google that statement.

OpenStudy (katecc379):

is this answer correct?

OpenStudy (katecc379):

i googled it it is true

OpenStudy (retireed):

according to Google... • The domain of a sequence consists of the counting numbers 1, 2, 3, 4, ... and the range consists of the terms of the sequence. Your statement is correct, so I think you are right.

OpenStudy (katecc379):

Thanks. Can you check the other attachment i sent? i know the sum is 168 I'm just stuck

OpenStudy (retireed):

I think the answer is going to be either c or d, not b let check again

OpenStudy (katecc379):

well the sum is 168 i know that, so then it must be d huh?

OpenStudy (retireed):

if it is 168 then it must be d since when i=1 a1 = 42

OpenStudy (sshayer):

it is last option

OpenStudy (sshayer):

\[a _{n}=ar ^{n-1}\] \[\sum_{i=1}^{\infty}a _{i}=\sum_{i=1}^{\infty}42\left( \frac{ 3 }{ 4 } \right)^{i-1}=\frac{ 42 }{ 1-\frac{ 3 }{ 4 } }=?\]

OpenStudy (sshayer):

for first sequence is first term,second term, third term,... so n=1,2,3,...

OpenStudy (sshayer):

for second \[common~ratio~r=\frac{ 3 }{ 4 }<1\] \[s _{\infty}=\frac{ a }{ 1-r }\]